The bright light emitted by sodium lamps comes from a transition between electrons at an energy level of – 3.04 eV falling to an energy level of -5.14 eV. Calculate the wavelength of the light emitted


1
Expert's answer
2021-01-19T07:09:02-0500

The wavelenght "\lambda" of the light is connected with its energy "E" in the following way:


"E = \dfrac{hc}{\lambda}"

where "h = 6.58\times 10^{-16}eV\cdot s" is the Plank's constant, and "c = 3\times 10^{8} m/s" is the speed of light.

On the other hand, the energy of the emitted light is equal to the difference between electon's energy levels:


"E = |-5.14eV - 3.04eV| = 8.18eV"

Expressing "\lambda", obtain:


"\lambda = \dfrac{hc}{E}\\
\lambda = \dfrac{6.58\times 10^{-16}\cdot 3\times 10^8}{8.18} \approx 2.41\times 10^{-8}m"

Answer. "2.41\times 10^{-8}m".